Consider an all-equity firm with 50,000 shares outstanding and a required rate of return of 12%/year.

PartA - What is the total value of the firm if it expects to earn $25,000/year in perpetuity, starting in 1 year from today, if it does not take on any new projects.

PartB - The firm has an unusual investment opportunity, which requires it to invest $10,000 today, and makes 9 more annual investments, at the end of each year for 9 more years, each investment decreasing by 15% from the previous year. In return, the investments expect to generate an additional $15,000 at the end of the 10th year, with additional annual cash flows growing after that, in perpetuity at 5%/year. What is the value of the investment opportunity?

PartC - What is the value of the firm, and the price per share of the firm, with the investment opportunity?
Answer #1

1
=25000/12%=208333.333333333

2
=10000*(1-(0.85/1.12)^10)/(1-(0.85/1.12))+15000/1.12^9*1.05/(12%-5%)=119989.296959095

3.
Value of the firm=(208333.333333333+119989.296959095)=328322.6303

Price per share=328322.6303/50000=6.566452606